A painter can finish painting a house in 5 hours. Her assistant takes 7 hours to finish the same job. How long would it take for them to complete the job if they were working together?
Answer by Alan3354(69443)

1 does 1/5 of the job per hour
The other does 1/7 per hour.
Together, 1/5 + 1/7 = 12/35 per hour
--> 35/12 hour/job
-----------------------
Use the shortcut, product over sum
5*7/(5+7) = 35/12